OpenStudy (armyrangers):

A weather balloon holds 2,850 cubic meters of helium. The density of helium is 0.1685 kilograms per cubic meter. How many kilograms of helium does the balloon contain? Please round your answer to the nearest whole number. Only enter a numerical value in the answer blank.

OpenStudy (anonymous):

480.225

OpenStudy (armyrangers):

thanks

OpenStudy (anonymous):

NP Glad to help :D

OpenStudy (accessdenied):

I agree with answer, but we should not just give the answer if we can. Knowing how to convert units is a huge skill in most computational Science! Unit analysis: \(\text{Volume}\times \text{Density} = \cancel{ \text{cubic meters}} \times \dfrac{\text{kilograms}}{\cancel{\text{cubic meters}}} = \text{kilograms}\) The cubic meters cancel like any numbers in both numerator and denominator, leaving kilograms. 2,850 [[cubic meters]], times 0.1685 [[kilograms per cubic meter]] = 480.225 [[kilograms]].

OpenStudy (anonymous):

:P